OpenStudy (anonymous):

simplify: 4t^2-1/ 15t+10 divided by 4t^2-4t+1/ 3t^2+5t+2 help please?

OpenStudy (anonymous):

\[\dfrac{\frac{4t^2-1}{15t+10}}{\frac{4t^2-4t+1}{3t^2+5t+2}}\]

OpenStudy (anonymous):

=\[\frac{4t^2-1}{15t+10}\frac{3t^2+5t+2}{4t^2-4t+1}\] =\[\frac{(2t-1)(2t+1)}{5(3t+2)}\frac{(3t+2)(t+1)}{(2t-1)^2}\] =\[\frac{(2t+1)(t+1)}{5(2t-1)}\]